# Exploring System 1 and 2 communication for latent reasoning in LLMs

### v1

@@ Abstract
Should LLM reasoning live in a separate module, or within a single model’s forward pass and representational space? We study dual-architecture latent reasoning, where a fluent Base exchanges latent messages with a Coprocessor, and test two hypotheses aimed at improving latent communication over Liu et al. (2024b): (H1) increase channel capacity; (H2) learn communication via joint finetuning. Under matched latent-token budgets on GPT-2 and Qwen-3, H2 is consistently strongest while H1 yields modest gains. A unified soft-embedding baseline—a single model with the same forward pass and shared representations, using the same latent-token budget—nearly matches H2 and surpasses H1, suggesting current dual designs mostly add compute rather than qualitatively improving reasoning. Across GSM8K, ProsQA, and a Countdown stress test with increasing branching factor, scaling the latent-token budget beyond small values fails to improve robustness. Latent analyses show overlapping subspaces with limited specialization, consistent with weak reasoning gains. We conclude dual-model latent reasoning remains promising in principle, but likely requires objectives and communication mechanisms that explicitly shape latent spaces for algorithmic planning.
@@ 1 Introduction
Large language models (LLMs) trained with web-scale pretraining and alignment have achieved impressive zero-shot reasoning capabilities across diverse tasks (Dubey et al., 2024; Hurst et al., 2024; Yang et al., 2025; Liu et al., 2024a). Despite their strong performance, LLMs are often seen as “fast and fluent” rather than genuinely deliberative — resembling the intuitive, System-1 side of dual-process theories of cognition (Kahneman, 2011). Indeed, recent surveys explicitly describe progress in reasoning LLMs as a shift from System-1-like heuristics to System-2-style deliberation (Li et al., 2025b), highlighting the need for architectures that support more structured reasoning.
The dominant approach today is chain-of-thought (CoT) reasoning (Wei et al., 2022; Liu et al., 2024a), where intermediate steps are verbalized in natural language. While effective, CoT incurs substantial token-level overhead, limits abstraction bandwidth, and constrains inference unnecessarily to the sequential, symbolic space of text (Chen et al., 2025; Qu et al., 2025). As model sizes and context window grow, these inefficiencies become increasingly prohibitive, motivating the search for more compact and expressive reasoning representations.
Latent reasoning (Hao et al., 2024; Liu et al., 2024b; Geiping et al., 2025) offers an alternative that enables the model to perform multi-step inference internally within its continuous hidden states, surfacing only the final answer. This paradigm promises two key advantages. First, reasoning in high-dimensional embeddings provides vastly greater expressive bandwidth than token sequences, potentially allowing richer intermediate computations (Zhang et al., 2025). Second, for combinatorial problems, operating over structured latent abstractions can dramatically reduce the effective search space (Geiping et al., 2025). Such ideas find parallels in cognitive science, where humans are believed to reason in internal “mentalese" before translating thoughts into language (Fodor, 1975).
Most latent-reasoning methods still ask a single network to do both fast association and slow deliberation; e.g., Coconut (Hao et al., 2024) feeds the model’s last hidden state back as input, forcing the same space to support next-token prediction and a putative “language of thought,” creating a representational tug-of-war. Neuroscience evidence instead points to partially distinct substrates (PFC for deliberative control; striatal circuits for habitual responses (Miller and Cohen, 2001; O’Reilly and Frank, 2006; Dolan and Dayan, 2013)), suggesting that separating roles could help. KV-cache Coprocessors (Liu et al., 2024b) fit this separation, but reported gains are limited; we argue the bottleneck is latent communication between modules.
@@
    We introduce two communication-oriented variants of Liu et al. (2024b)’s dual architecture and show better performance under matched latent-token budgets.
@@ 2 Related work
Reasoning in latent space. Latent-space methods shift multi-step inference from tokenized CoT into a small set of latent variables, decoupling compute from text length. Representative approaches include Coconut, which feeds continuous “thoughts” back into the model (Hao et al., 2024); differentiable KV-cache augmentation, which separates a Base token predictor from a Coprocessor that edits the cache (Liu et al., 2024b); and compressed/implicit CoT that learns dense contemplation tokens or plans without emitting text (Cheng et al., 2024; Kurtz et al., 2024). Recent surveys synthesize this trend and its claimed benefits (Li et al., 2025a; Zhu et al., 2025). Despite clear token-efficiency gains, reported improvements on reasoning are mixed, motivating our study of how to train dual-model systems for effective latent communication.
Communication and coordination between models. Our focus on “latent communication” between a Base and a Coprocessor connects to two strands. First, teacher–student transfer suggests that sharing an initialization can facilitate implicit trait transfer, even without explicit supervision (Cloud et al., 2025); this supports initializing the Coprocessor from the Base to align internal representations. Second, multi-agent prompting frameworks use inter-model dialogue to improve reliability (Du et al., 2023; Liang et al., 2023), though recent analyses find gains can be brittle or dataset-dependent (Wang et al., 2024). Finally, our unified soft-embedding baseline is grounded in continuous-prompt methods that endow a single network with extra latent capacity via trainable prefix/prompt vectors (Lester et al., 2021; Li and Liang, 2021; Goyal et al., 2023), providing a strong, parameter-efficient alternative to dual-model designs.
@@ 3.1 Problem setting
Figure 1: Overview of the deliberation–in–KV-cache architecture of Liu et al. (2024b) and our two variants designed to strengthen cross-module communication.
Setup and notation. Let $B_{\theta}$ be a frozen, pretrained LLM (Base) with parameters $\theta$. Given a prompt $x$ and target $y$, a forward pass of $B_{\theta}$ produces per-layer key–value caches $\{(K_{\ell}(x;\theta),V_{\ell}(x;\theta))\}_{\ell=1}^{L}$, abbreviated $KV_{\theta}(x)$. A Coprocessor $C_{\phi}$ with parameters $\phi$ reads $KV_{\theta}(x)$ together with $N_{L}$ learnable soft tokens and emits a latent sequence $Z\in\mathbb{R}^{N_{L}\times d}$. At decoding time, $Z$ is injected back into $B_{\theta}$ (in a manner that depends on the variant below), after which the Base generates $y$.
@@
where $Z=C_{\phi}\big(KV_{\theta}(x),N_{L}\text{-soft}\big)$ and $\operatorname{inject}(\cdot)$ denotes the chosen mechanism for feeding $Z$ (or the Coprocessor’s cache) back to the Base.
Pipeline (as in Liu et al. (2024b)). We follow the standard three-stage process (cf. Fig. 1, left):
@@
    Latent augmentation. Run $C_{\phi}$ on $KV_{\theta}(x)$ plus $N_{L}$ learnable soft tokens to produce a latent sequence $Z$. In Liu et al. (2024b), only the final layer’s hidden states of $C_{\phi}$ are used as input embeddings to $B_{\theta}$.
@@ 3.2 Experimental variants
i.e., concatenation along the sequence (cache) dimension. [^1] We denote sequence-axis concatenation by $\cdot\,$$\cdot$]. Optimization: $\theta$ remains frozen; only $\phi$ and the soft tokens are trained. Motivation: With a frozen Base, steering only via input embeddings gives the Coprocessor influence at the first hidden layer. Cache-level augmentation propagates the latent signal through all layers, potentially enabling richer, layer-wise “latent communication”. Prediction: With $\theta$ frozen, cache augmentation will outperform embedding-only feedback (Liu et al., 2024b) at matched latent-token budgets $N_{L}$.
Hypothesis 2 — Co-finetuned dual-model. Change: Same injection mechanism as the reference system in Liu et al. (2024b) (latents as input embeddings to the Base), but we unfreeze the Base. Optimization: Update both $\theta$ and $\phi$ jointly (plus soft tokens) under the log-likelihood objective above. Motivation: Joint training allows $B_{\theta}$ to learn to “listen” to the Coprocessor’s messages rather than treating them as exogenous noise. Although this increases the number of trainable parameters, it isolates whether improved cross-module interaction—rather than mere extra compute—drives downstream gains. Prediction: For the same latent injection as Liu et al. (2024b), co-finetuning will outperform the frozen-Base counterpart at matched $N_{L}$, reflecting learned communication rather than added compute.
@@ 4.1 Large scale data training
We replicate the deliberation KV-cache pipeline of Liu et al. (2024b), but adapt it to a tighter compute budget and to two model families.
Base models: GPT-2 (124 M) (Radford et al., 2019) and Qwen-3 (0.6 B) (Yang et al., 2025) replace the 2-B-parameter Gemma used in the Liu et al. (2024b) paper. This choice allows us to complete all runs within a 3-day window on an 8 × A100 (80 GB) node.
@@
Dataset. All large-scale training uses FineWeb-Edu-100BT (Lozhkov et al., 2024).
Sequence & latent parameters: We use sequence length S=1024S{=}1024, latent augmentations per sequence M=64M{=}64, ahead tokens $N_{A}{=}16$ for back-propagating the loss into the Coprocessor (Liu et al. (2024b) use $2048/128/16$), and $N_{L}{=}16$ latents. Figure 2 C/D show that validation perplexity decreases almost linearly as $N_{L}$ grows under Hypothesis 2. Training cost scales with the effective per-example context $S+M\!\cdot\!(N_{L}{+}N_{A})$, so larger $N_{L}$ substantially raises wall-clock. Within our budget, $N_{L}{=}16$ provided a practical operating point, while Liu et al. (2024b) report their strongest results around $N_{L}{=}32$ on larger bases.
Three-pass training loop implementation (Fig. 5): Liu et al. (2024b) describe “an efficient training framework … in one forward pass”, enabled by a custom attention mask that scales to large datasets. Such a mask must support multiple augmentations (M) per sequence; otherwise only one augmentation is trained, yielding two orders of magnitude less signal per token. However, a single forward pass could allow the Base model to shortcut the intended latent computation—performing both next-token prediction and latent augmentation itself, thereby collapsing the two representational spaces and defeating the purpose of the technique. We cannot verify whether Liu et al. (2024b) avoid this, as their code is not public; the attention masks they describe imply that only one model processes the pass, but the exact implementation remains unclear. To keep our dataset large while preventing this shortcut, we adopt a strict three-pass loop (Figure 5) whose attention masks still allow M augmentations of each sequence.
@@ 4.1.1 Baselines
Baseline 1 - Base model + continued pretraining: Liu et al. (2024b), compare their dual-model to the initial base checkpoint, meaning the dual system has seen far more data. Instead, we keep the architecture unchanged and simply continue pretraining the base model on the same number of tokens used by our dual-model runs.
Baseline 2 - Base model + $N_{L}$ soft embeddings: To test whether a single network can absorb the “System-2” role, we continue pretraining the base model alone while attaching the same number of new $N_{L}$ learnable soft tokens used by the Coprocessor setup (similar in spirit to Goyal et al. (2023), but with untied, slot-specific embeddings rather than a single shared token). This contrasts a unified versus dual-network design under identical data and latent embedding budgets. See App. 6 for visualization.
@@ 4.1.2 Results
Qualitative replication of Liu et al. (2024b). Substituting the 2-B parameter Gemma with much smaller GPT-2 (124 M) and Qwen-3 (0.6 B) still produces the qualitative trends reported by Liu et al. (2024b)(Fig. 2A/B). Perplexity (ppl) drops by $\simeq$ 2 for GPT-2 and $\simeq$ 2.5 for Qwen-3—an order-of-magnitude larger reduction than that reported in the original paper setup. On GPT-2 the dual-model also raises mean benchmark accuracy by +2.0 percentage point ($p$) (Table 1). For Qwen-3, however, lower perplexity does not translate into higher benchmark scores; continued pretraining alone even degrades accuracy. Given Qwen-3’s heavy post-training, such brittleness under distribution shift is expected. We therefore report all numbers relative to the continued-pretraining baseline.
Both hypotheses outperform Liu et al. (2024b). Relative to Liu et al.’s dual-model, Hypothesis 1 (frozen Base, cache concatenation) improves average accuracy by +1.5 $p$ on GPT-2 and +4.4 $p$ on Qwen-3. Hypothesis 2 (co-finetuned) improves by +3.7 $p$ (GPT-2) and +6.6 $p$ (Qwen-3). In ppl, the variants are $\approx 2$ and $\approx 4$ lower, respectively (Fig. 2A/B).
Liu model versus data-matched continued pretraining. Relative to our data-matched Baseline 1 (same number of training tokens), the Liu et al.’s dual-model exhibits higher ppl ($\approx{+}1.5$) and mixed accuracy: 2.7-2.7 $p$ on GPT-2 but +2.1+2.1 $p$ on Qwen-3. Averaged across families the net change is $\approx{-}0.3$ $p$, underscoring the value of stricter baselines.
@@
  • GPT-2 benchmarks: Averaged over the five tasks in Table 1, Hyp. 1 trails by 1.8-1.8 $p$ (46.3 vs. 48.1%), while Hyp. 2 is only +0.4+0.4 $p$ higher (48.5%).
  • Qwen-3 benchmarks: Dual models fare better: Hyp. 1 +6.5+6.5 $p$ (47.2 vs. 40.7%); Hyp. 2 +8.7+8.7 $p$ (49.4%).
Summary and caveat. At first glance both dual-model variants look successful: they outperform the Liu et al. (2024b) baseline and, on Qwen-3, yield sizeable benchmark gains. But the picture shifts once we introduce the soft-embedding control. That unified model matches the dual systems in token budget and latent capacity while using only half as many trainable weights, yet it equals or exceeds Hyp. 1 and comes within a hair of Hyp. 2. This pattern indicates that the Coprocessor is not just “adding compute”—it is adding it inefficiently. A single LLM with the same aggregate parameter count would almost certainly do better. Accordingly, the next section turns to reasoning-specific benchmarks to see whether the dual architecture offers any benefit that a parameter-matched, soft-prompted model cannot already provide.

### v2

@@ Abstract
Should LLM reasoning live in a separate module, or within a single model’s forward pass and representational space? We study dual-architecture latent reasoning, where a fluent Base exchanges latent messages with a Coprocessor, and test two hypotheses aimed at improving latent communication over liu2024deliberationlatentspacedifferentiable: (H1) increase channel capacity; (H2) learn communication via joint finetuning. Under matched latent-token budgets on GPT-2 and Qwen-3, H2 is consistently strongest while H1 yields modest gains. A unified soft-embedding baseline—a single model with the same forward pass and shared representations, using the same latent-token budget—nearly matches H2 and surpasses H1, suggesting current dual designs mostly add compute rather than qualitatively improving reasoning. Across GSM8K, ProsQA, and a Countdown stress test with increasing branching factor, scaling the latent-token budget beyond small values fails to improve robustness. Latent analyses show overlapping subspaces with limited specialization, consistent with weak reasoning gains. We conclude dual-model latent reasoning remains promising in principle, but likely requires objectives or training schedules that explicitly shape latent spaces for algorithmic planning.
@@ 1 Introduction
Large language models (LLMs) trained with web-scale pretraining and alignment have achieved impressive zero-shot reasoning capabilities across diverse tasks (dubey2024llama; hurst2024gpt; yang2025qwen3; liu2024deepseek). Despite their strong performance, LLMs are often seen as “fast and fluent” rather than genuinely deliberative — resembling the intuitive, System-1 side of dual-process theories of cognition (kahneman2011thinking). Indeed, recent surveys explicitly describe progress in reasoning LLMs as a shift from System-1-like heuristics to System-2-style deliberation (li202512surveyreasoning), highlighting the need for architectures that support more structured reasoning.
The dominant approach today is chain-of-thought (CoT) reasoning (wei2022chain; liu2024deepseek), where intermediate steps are verbalized in natural language. While effective, CoT incurs substantial token-level overhead, limits abstraction bandwidth, and constrains inference unnecessarily to the sequential, symbolic space of text (chen2025towards; qu2025survey). As model sizes and context window grow, these inefficiencies become increasingly prohibitive, motivating the search for more compact and expressive reasoning representations.
Latent reasoning (coconut; liu2024deliberationlatentspacedifferentiable; geiping2025scaling) offers an alternative that enables the model to perform multi-step inference internally within its continuous hidden states, surfacing only the final answer. This paradigm promises two key advantages. First, reasoning in high-dimensional embeddings provides vastly greater expressive bandwidth than token sequences, potentially allowing richer intermediate computations (zhang2025soft). Second, for combinatorial problems, operating over structured latent abstractions can dramatically reduce the effective search space (geiping2025scaling). Such ideas find parallels in cognitive science, where humans are believed to reason in internal “mentalese" before translating thoughts into language (fodor1975language).
Most latent-reasoning methods still ask a single network to do both fast association and slow deliberation; e.g., Coconut (coconut) feeds the model’s last hidden state back as input, forcing the same space to support next-token prediction and a putative “language of thought,” creating a representational tug-of-war. Neuroscience evidence instead points to partially distinct substrates (PFC for deliberative control; striatal circuits for habitual responses (MillerCohen2001; OReillyFrank2006; DolanDayan2013)), suggesting that separating roles could help. KV-cache Coprocessors (liu2024deliberationlatentspacedifferentiable) fit this separation, but reported gains are limited; we argue the bottleneck is latent communication between modules.
@@
    We introduce two communication-oriented variants of liu2024deliberationlatentspacedifferentiable’s dual architecture and show better performance under matched latent-token budgets.
@@
  3. 3.
    We validate this hypothesis by demonstrating that explicit regularization can force latent specialization and restore reasoning signatures on combinatorial tasks, while highlighting the inherent trade-off this creates with general language modeling.
@@ 2 Related work
Reasoning in latent space. Latent-space methods shift multi-step inference from tokenized CoT into a small set of latent variables, decoupling compute from text length. Representative approaches include Coconut, which feeds continuous “thoughts” back into the model (coconut); differentiable KV-cache augmentation, which separates a Base token predictor from a Coprocessor that edits the cache (liu2024deliberationlatentspacedifferentiable); and compressed/implicit CoT that learns dense contemplation tokens or plans without emitting text (cheng2024ccot; kurtz2024quietstar). Recent surveys synthesize this trend and its claimed benefits (li2025latentcotsurvey; zhu2025latentcot). Despite clear token-efficiency gains, reported improvements on reasoning are mixed, motivating our study of how to train dual-model systems for effective latent communication.
Communication and coordination between models. Our focus on “latent communication” between a Base and a Coprocessor connects to two strands. First, teacher–student transfer suggests that sharing an initialization can facilitate implicit trait transfer, even without explicit supervision (cloud2025subliminal); this supports initializing the Coprocessor from the Base to align internal representations. Second, multi-agent prompting frameworks use inter-model dialogue to improve reliability (du2023mad; liang2023mad), though recent analyses find gains can be brittle or dataset-dependent (wang2024aremultiagent). Finally, our unified soft-embedding baseline is grounded in continuous-prompt methods that endow a single network with extra latent capacity via trainable prefix/prompt vectors (lester2021prompttuning; li2021prefixtuning; goyalthink), providing a strong, parameter-efficient alternative to dual-model designs.
@@ 3.1 Problem setting
Figure 1: Overview of the deliberation–in–KV-cache architecture of liu2024deliberationlatentspacedifferentiable and our two variants designed to strengthen cross-module communication.
Setup and notation.
Let $B_{\theta}$ be a frozen, pretrained LLM (Base) with parameters $\theta$. Given a prompt $x$ and target $y$, a forward pass of $B_{\theta}$ produces per-layer key–value caches $\{(K_{\ell}(x;\theta),V_{\ell}(x;\theta))\}_{\ell=1}^{L}$, abbreviated $KV_{\theta}(x)$. We define a set of $M$ augmentation sites indices $\mathcal{T}=\{t_{1},\dots,t_{M}\}$ within the sequence $x$. At each site $t_{m}$, a Coprocessor $C_{\phi}$ with parameters $\phi$ reads $KV_{\theta}(x)$ up to $t_{m}$ together with $N_{L}$ learnable soft tokens and emits a latent sequence $Z\in\mathbb{R}^{N_{L}\times d}$. At decoding time, $Z$ is injected back into $B_{\theta}$ to predict the subsequent tokens. During training, gradients flow from the next $N_{A}$ “ahead” tokens (the supervision window) back to the latents.
@@
where $Z=C_{\phi}\big(KV_{\theta}(x),N_{L}\text{-soft}\big)$ and $\operatorname{inject}(\cdot)$ denotes the chosen mechanism for feeding $Z$ (or the Coprocessor’s cache) back to the Base. See App 6.1 for a more detailed formulation.
Pipeline (as in liu2024deliberationlatentspacedifferentiable). We follow the standard three-stage process (cf. Fig. 1, left):
@@
    Latent augmentation. Run $C_{\phi}$ on $KV_{\theta}(x)$ plus $N_{L}$ learnable soft tokens to produce a latent sequence $Z$. In liu2024deliberationlatentspacedifferentiable, only the final layer’s hidden states of $C_{\phi}$ are used as input embeddings to $B_{\theta}$.
@@ 3.2 Experimental variants
i.e., concatenation along the sequence (cache) dimension. [^1] We denote sequence-axis concatenation by $\cdot\,$$\cdot$]. Optimization: $\theta$ remains frozen; only $\phi$ and the soft tokens are trained. Motivation: With a frozen Base, steering only via input embeddings gives the Coprocessor influence at the first hidden layer. Cache-level augmentation propagates the latent signal through all layers, potentially enabling richer, layer-wise “latent communication”. Prediction: With $\theta$ frozen, cache augmentation will outperform embedding-only feedback (liu2024deliberationlatentspacedifferentiable) at matched latent-token budgets $N_{L}$.
Hypothesis 2 — Co-finetuned dual-model. Change: Same injection mechanism as the reference system in liu2024deliberationlatentspacedifferentiable (latents as input embeddings to the Base), but we unfreeze the Base. Optimization: Update both $\theta$ and $\phi$ jointly (plus soft tokens) under the log-likelihood objective above. Motivation: Joint training allows $B_{\theta}$ to learn to “listen” to the Coprocessor’s messages rather than treating them as exogenous noise. Although this increases the number of trainable parameters, it isolates whether improved cross-module interaction—rather than mere extra compute—drives downstream gains. Prediction: For the same latent injection as liu2024deliberationlatentspacedifferentiable, co-finetuning will outperform the frozen-Base counterpart at matched $N_{L}$, reflecting learned communication rather than added compute.
@@ 4.1 Large scale data training
We replicate the deliberation KV-cache pipeline of liu2024deliberationlatentspacedifferentiable, but adapt it to a tighter compute budget and to two model families.
Base models: GPT-2 (124 M) (radford2019language) and Qwen-3 (0.6 B) (yang2025qwen3) replace the 2-B-parameter Gemma used in the liu2024deliberationlatentspacedifferentiable paper. This choice allows us to complete all runs within a 3-day window on an 8 × A100 (80 GB) node.
@@
Dataset. All large-scale training uses FineWeb-Edu-100BT (lozhkov2024fineweb-edu).
Sequence & latent parameters: We use sequence length S=1024S{=}1024, latent augmentations per sequence M=64M{=}64, ahead tokens $N_{A}{=}16$ for back-propagating the loss into the Coprocessor (liu2024deliberationlatentspacedifferentiable use $2048/128/16$), and $N_{L}{=}16$ latents. Figure 2 C/D show that validation perplexity decreases almost linearly as $N_{L}$ grows under Hypothesis 2. Training cost scales with the effective per-example context $S+M\!\cdot\!(N_{L}{+}N_{A})$, so larger $N_{L}$ substantially raises wall-clock. Within our budget, $N_{L}{=}16$ provided a practical operating point, while liu2024deliberationlatentspacedifferentiable report their strongest results around $N_{L}{=}32$ on larger bases. We provide further ablations of these hyper-parameters in App 11
Three-pass training loop implementation (Fig. 7): liu2024deliberationlatentspacedifferentiable describe “an efficient training framework … in one forward pass”, enabled by a custom attention mask that scales to large datasets. Such a mask must support multiple augmentations (M) per sequence; otherwise only one augmentation is trained, yielding two orders of magnitude less signal per token. However, we hypothesize that a single forward pass creates an optimization shortcut: since the Base model is pre-trained and competent while the Coprocessor begins as initialized noise, the joint system is incentivized to ignore the noisy latent channel in favor of the Base’s internal representations. This reflects the ideas around path of least resistance in deep learning (bowman2015generating; geirhos2020shortcut). Furthermore, it is technically unclear how to implement such a bidirectional dependency (Base $\to$ Coprocessor $\to$ Base) within a single forward pass, and liu2024deliberationlatentspacedifferentiable do not provide implementation details to clarify this. To ensure the Base strictly conditions on the Coprocessor’s output and to avoid optimization collapse, we adopt a strict three-pass loop (Figure 7) that architecturally enforces the dependency.
@@ 4.1.1 Baselines
Baseline 1 - Base model + continued pretraining: liu2024deliberationlatentspacedifferentiable, compare their dual-model to the initial base checkpoint, meaning the dual system has seen far more data. Instead, we keep the architecture unchanged and simply continue pretraining the base model on the same number of tokens used by our dual-model runs.
Baseline 2 - Base model + $N_{L}$ soft embeddings: To test whether a single network can absorb the “System-2” role, we continue pretraining the base model alone while attaching the same number of new $N_{L}$ learnable soft tokens used by the Coprocessor setup (similar in spirit to goyalthink, but with untied, slot-specific embeddings rather than a single shared token). This contrasts a unified versus dual-network design under identical data and latent embedding budgets. See App. 6 for visualization.
@@ 4.1.2 Results
Qualitative replication of liu2024deliberationlatentspacedifferentiable. Substituting the 2-B parameter Gemma with much smaller GPT-2 (124 M) and Qwen-3 (0.6 B) still produces the qualitative trends reported by liu2024deliberationlatentspacedifferentiable(Fig. 2A/B). Perplexity (ppl) drops by $\simeq$ 2 for GPT-2 and $\simeq$ 2.5 for Qwen-3—an order-of-magnitude larger reduction than that reported in the original paper setup. On GPT-2 the dual-model also raises mean benchmark accuracy by +2.0 percentage point (pp) (Table 1). For Qwen-3, however, lower perplexity does not translate into higher benchmark scores; continued pretraining alone even degrades accuracy. Given Qwen-3’s heavy post-training, such brittleness under distribution shift is expected. We therefore report all numbers relative to the continued-pretraining baseline.
Both hypotheses outperform liu2024deliberationlatentspacedifferentiable. Relative to Liu et al.’s dual-model, Hypothesis 1 (frozen Base, cache concatenation) improves average accuracy by +1.5 pp on GPT-2 and +4.4 pp on Qwen-3. Hypothesis 2 (co-finetuned) improves by +3.7 pp (GPT-2) and +6.6 pp (Qwen-3). In ppl, the variants are $\approx 2$ and $\approx 4$ lower, respectively (Fig. 2A/B).
Liu model versus data-matched continued pretraining. Relative to our data-matched Baseline 1 (same number of training tokens), the Liu et al.’s dual-model exhibits higher ppl ($\approx{+}1.5$) and mixed accuracy: 2.7-2.7 pp on GPT-2 but +2.1+2.1 pp on Qwen-3. Averaged across families the net change is $\approx{-}0.3$ pp, underscoring the value of stricter baselines.
@@
  • GPT-2 benchmarks: Averaged over the five tasks in Table 1, Hyp. 1 trails by 1.8-1.8 pp (46.3 vs. 48.1%), while Hyp. 2 is only +0.4+0.4 pp higher (48.5%).
  • Qwen-3 benchmarks: Dual models fare better: Hyp. 1 +6.5+6.5 pp (47.2 vs. 40.7%); Hyp. 2 +8.7+8.7 pp (49.4%).
Summary and caveat. At first glance both dual-model variants look successful: they outperform the liu2024deliberationlatentspacedifferentiable baseline and, on Qwen-3, yield sizeable benchmark gains. But the picture shifts once we introduce the soft-embedding control. That unified model matches the dual systems in token budget and latent capacity while using only half as many trainable weights, yet it equals or exceeds Hyp. 1 and comes within a hair of Hyp. 2. This pattern indicates that the Coprocessor is not just “adding compute”—it is adding it inefficiently. A single LLM with the same aggregate parameter count indeed performs better, as we explicitly demonstrate in Appendix 10. Accordingly, the next section turns to reasoning-specific benchmarks to see whether the dual architecture offers any benefit that a parameter-matched, soft-prompted model cannot already provide.